Solve for x. The equation is 5/4x + 1/8x = 3/8 + x. Thank you for your help!
Answer by ewatrrr(24785)


Hi
Solving the following for x:
5/4x + 1/8x = 3/8 + x
10/8x + 1/8x = 3/8 + x
11/8x - x = 3/8
3/8x = 3/8
x = 1